# On an Optimal Stopping Problem with a Discontinuous Reward

### v3

@@ 4.4. Characterization of the Surrender Region
The statements of Proposition 4.6 (ii) and 4.7 (i) do not cover the case where $L(t)=0$. Heuristically, when $L(t)=0$ for all $0\leq t\leq T$, the discounted surrender value process $\{Y_{t}\}_{0\leq t\leq T}$, with $Y_{t}=e^{-rt}F_{t}g(t)$, is a martingale, so one might expect that all stopping times $\tau$ such that $0\leq\tau\leq T$ are optimal, implying that ${\mathcal{S}=[0,T)\times\mathbb{R}_{+}}$, see for instance Björk (2009), Proposition 21.2. However, this is incorrect; due to the time discontinuity of the reward function in (7), the discounted reward process $\{Z_{t}\}_{0\leq t\leq T}$, with $Z_{t}=e^{-rt}\varphi(t,F_{t})$, is a submartingale (see Lemma 3.1). Hence, the policyholder can always profit (on average) from holding on to the contract because of the guaranteed amount at maturity. The optimal stopping time is then $T$ and the surrender region is empty. This illustrates another major difference between the optimal stopping problem in (7) and the one involved in the pricing of standard American options.
The specificity of the VA pricing problem discussed above is also illustrated in the continuation premium representation in (28). Indeed, when ${L(t)=0}$ for all $t\in[0,T]$, then for any $x\in\mathbb{R}_{+}$, ${f(t,x)=\mathbb{E}[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}]>0}$. That is, the continuation premium $f$ is equal to the expected present value of the financial guarantee. Hence, $v(t,x)=\varphi(t,x)+f(t,x)>\varphi(t,x)$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, which implies that $\mathcal{S}=\emptyset$.
Propositions 4.6 and 4.7 can now be used to prove Theorem 4.3.
@@
Finally, since $x\in\mathcal{C}_{t}$, we have $v(t,\,x)>g(t)x$, so that $v(t,y)>g(t)y$, which implies $y\in\mathcal{C}_{t}$.
@@
Hence, $\mathcal{S}=\{(t,x)\in[0,T)\times\mathbb{R}_{+}|v(t,x)=\varphi(t,x)\}=\emptyset$, so that ${\tau_{t}^{x}=\inf\{t\leq s\leq T|v(s,F_{s}^{t,x})=\varphi(s,F_{s}^{t,x})\}=T}$. Now since $\tau_{t}^{x}$ is the smallest optimal stopping time for (7), as per Theorem 3.3 (iii), we conclude that it is unique (since all other optimal stopping times must be greater than $\tau_{t}^{x}$ and smaller than $T$).
@@ 4.5. Equivalence of the Optimal Stopping Problems
If $L(t)<0$, ${v(t,x)>h(t,x)}$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, where ${h(t,x)=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right]}$.
@@
To show that $\mathcal{S}\supseteq\widetilde{\mathcal{S}}$, we show $\mathcal{C}\subseteq\widetilde{\mathcal{C}}$. Fix $(t,x)\in\mathcal{C}$, so that ${v(t,x)>\varphi(t,x)}$. Since $L(t)<0$ for all $t\in[0,T)$, we have by Lemma 4.5 that ${v(t,x)>h(t,x)}$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. Thus, $v(t,x)>\max(\varphi(t,x),h(t,x))$, which confirms that $(t,x)\in\widetilde{\mathcal{C}}$. ∎
@@
$\begin{array}[]{cc}\tau^{x}_{t}&=\inf\{t\leq s\leq T|(s,F_{s}^{t,x})\in\mathcal{S}\},\\ \widetilde{\tau}^{x}_{t}&=\inf\{t\leq s\leq T|(s,F_{s}^{t,x})\in\widetilde{\mathcal{S}}\}.\end{array}$
@@
To provide some intuition for this result, we remark that the condition $L(t)<0$ is associated with a non-empty surrender region (see Proposition 4.6); at any point during the life of the contract, there is always a possibility that early surrender will become optimal in the future. Th, enthe value of the VA contract is strictly above that of the EPV of the maturity benefit, and if surrender occurs it will be due to the contract value being equal to the surrender benefit. When this is the case, all three stopping times coincide (see the discussion after the proof of Lemma 3.3).
@@ 5. Numerical Examples
In this section, we present two simple examples to highlight the specific features of the optimal stopping problem stemming from variable annuity pricing, when compared to those associated with the American call and put option pricing. In particular, with variable annuities, the surrender region can be disconnected and the optimal stopping boundary, discontinuous.
@@
$c_{1}(t)$ $=\frac{1}{100}\left(0.00889t^{2}-0.1330t+1\right),$
$c_{2}(t)$ $=\frac{1}{100}\left(-0.0333t+1\right),$
for $0\leq t\leq T$, with $T=15$. The surrender charge function is set to $g(t)=e^{-\kappa(T-t)}$, with $\kappa=0.0055$, which is a commonly used form in the variable annuity literature. The resulting surrender charge at inception of the contract is close to $8\%$, and it goes down to less than $0.5\%$ at $t=14$, one year before maturity.
The fee functions were chosen to illustrate some of the results obtained in Section 4. Function $c_{1}$ shows the impact of a change in the sign of $L(t)$; it is negative on $[0,5.17]\cup[9.79,15]$ and positive elsewhere. Function $c_{2}$ was chosen such that $L(t)>0$ for $t\in(13.5,T]$ for $c_{2}$, thus eliminating the risk-neutral surrender incentive in the last 1.5 year of the contract. In both cases, the annual fee rates vary between $0.5\%$ and $1.0\%$. In practice, the fee is typically set as a constant and the surrender charge only depends on time. Numerical examples in this setting are provided, for example, in Bernard et al. (2014b), MacKay (2014) and MacKay et al. (2017).
@@
Figure 1 shows the continuation regions (in red) associated to the optimal stopping problem in (7) for the two fee functions $c_{1}$ and $c_{2}$. The surrender regions (in white) are empty between years $5$ and $10$ for the fee function $c_{1}$, and between years $13.5$ and $15$ for $c_{2}$. This corresponds to the time intervals during which $L(t)>0$, illustrating a result of Proposition 4.7 (ii). Thus, the optimal surrender boundary is discontinuous when the fee is given by $c_{1}$. It is already known from the literature of variable annuity that the surrender region can have different shapes depending on the fee function, see for example MacKay et al. (2017), Figures 4 and 5. However, such disconnected sets for the surrender region have not been observed in prior numerical work on similar problems, see among others Bernard et al. (2014b), MacKay et al. (2017), Kang and Ziveyi (2018), and MacKay et al. (2023). This simple example illustrates important differences between the optimal stopping problem involved in variable annuities pricing and the one associated with American call and put options.
Another important feature highlighted in the second example (with fee function $c_{2}$) is that we no longer have $\lim_{t\rightarrow T}b(t)=G$. This is because the surrender region is empty for $t>13.5$, just before maturity (since $L(t)>0$). Moreover, in this particular case, since $\mathcal{S}\cap([13.5,T)\times\mathbb{R}_{+})=\emptyset$, we have $v(t,x)=h(t,x)$ for $(t,x)\in[13.5,T)$, so that $\widetilde{\mathcal{S}}\cap([13.5,T)\times\mathbb{R}_{+})=[13.5,T)\times\mathbb{R}_{+}$, providing another example in which the optimal stopping problems with the discontinuous and the continuous reward functions have different surrender (and continuation) regions.
Finally, we note that in our figures, the surrender region becomes empty slightly before $t=5$ for the fee function $c_{1}$, and before $t=13.5$ for $c_{2}$. This is likely due to the $y$-axis stopping at 500; it may be that there exists a surrender region at $t=5$ for account values that are (much) higher than 500, and that the surrender boundary is continuous in a general sense, “reaching infinity” continuously when the sign of $L$ changes. This would also mean that theoretically, we could show that $L(t)\leq 0$ if and only if $\mathcal{S}_{t}$ is non-empty. Verifying such properties of the surrender boundary is left for future research.
@@ 6. Conclusion
In particular, the continuous reward representation allows us to use existing results from American option pricing theory to obtain the continuity of the value function. We also prove its convexity in the state variable, its Lipschitz property in $x$ and characterize its continuity in $t$. Then, we derive an integral expression for the early surrender premium, generalizing the results of Bernard et al. (2014a) to general time-dependent fee and surrender charge functions. This second representation of the value function is also known as the early exercise premium representation in the American option pricing literature. From there, we develop a third representation for the value function in terms of the current surrender value and an integral expression that only takes value in the continuation region. We call this third representation the continuation premium representation.
@@
    $\leq\mathbb{E}\left[e^{-r(\tau^{x}_{t}-t)}\left|xe^{(r-\sigma^{2}/2)(\tau^{x}_{t}-t)-\int_{0}^{\tau_{t}^{x}-t}c(t+s)\,\mathrm{d}s+\sigma W_{\tau^{x}_{t}-t}}\right.\right.$
@@
    $\quad=\sup_{\widetilde{\tau}\in[0,1]}\mathbb{E}\left[e^{-r\widetilde{\tau}(T-t)}\varphi\left(t+\widetilde{\tau}(T-t),xe^{(r-\sigma^{2}/2)\widetilde{\tau}(T-t)+\int_{t}^{\widetilde{\tau}(t+T-t)}c(u)\,\,\mathrm{d}u+\sigma\sqrt{T-t}W_{\widetilde{\tau}}}\right)\right],$
@@
    $v(t,x)=\mathbb{E}\left[e^{-r\widetilde{\tau}_{t}(T-t)}\varphi(t+\widetilde{\tau}_{t}(T-t),\widetilde{F}^{t,x}_{t+\widetilde{\tau}(T-t))}\right].$
    Fix $0\leq s\leq t\leq T$. Since $v$ is non-monotone in $t$ (see Remark 4.5), we consider two cases: $v(s,x)\geq v(t,x)$ and $v(s,x)<v(t,x)$
@@
    $\qquad\qquad+|\sqrt{T-t}-\sqrt{T-s}|x\sigma e^{|r-\sigma^{2}/2|T+\int_{0}^{T}c(u)\,\mathrm{d}u}\mathbb{E}[M_{1}e^{\sigma\sqrt{T}M_{1}}]$
    $\quad=\widetilde{C}_{1}|t-s|+C_{2}|\sqrt{T-t}-\sqrt{T-s}|,$
@@
    $\quad=C_{1}|t-s|+C_{2}\left|\sqrt{T-s}-\sqrt{T-t}\right|$

### v4

@@ 4.4. Characterization of the Surrender Region
The statements of Proposition 4.5 (ii) and 4.6 (i) do not cover the case where $L(t)=0$. Heuristically, when $L(t)=0$ for all $0\leq t\leq T$, the discounted surrender value process $\{Y_{t}\}_{0\leq t\leq T}$, with $Y_{t}=e^{-rt}F_{t}g(t)$, is a martingale, so one might expect that all stopping times $\tau$ such that $0\leq\tau\leq T$ are optimal, implying that $\mathcal{S}=[0,T)\times\mathbb{R}_{+}$}, see for instance Björk (2009), Proposition 21.2. However, this is incorrect; due to the time discontinuity of the reward function in (7), the discounted reward process $\{Z_{t}\}_{0\leq t\leq T}$, with $Z_{t}=e^{-rt}\varphi(t,F_{t})$, is a submartingale (see Lemma 3.1). Hence, the policyholder can always profit (on average) from holding on to the contract because of the guaranteed amount at maturity. The optimal stopping time is then $T$ and the surrender region is empty. This illustrates another major difference between the optimal stopping problem in (7) and the one involved in the pricing of standard American options.
The specificity of the VA pricing problem discussed above is also illustrated in the continuation premium representation in (29). Indeed, when L(t)=0{L(t)=0} for all $t\in[0,T]$, then for any $x\in\mathbb{R}_{+}$, $f(t,x)=\mathbb{E}[e^{-r(T-t)}(G-F_{T}^{t,x})_{+}]>0$}. That is, the continuation premium $f$ is equal to the expected present value of the financial guarantee. Hence, $v(t,x)=\varphi(t,x)+f(t,x)>\varphi(t,x)$ for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, which implies that $\mathcal{S}=\emptyset$.
Propositions 4.5 and 4.6 can now be used to prove Theorem 4.3.
@@
Finally, since $x\in\mathcal{C}_{t}$, we have v(t,x)>g(t)xv(t,\,x)>g(t)x, so that $v(t,y)>g(t)y$, which implies $y\in\mathcal{C}_{t}$.
@@
Hence, $\mathcal{S}=\{(t,x)\in[0,T)\times\mathbb{R}_{+}|v(t,x)=\varphi(t,x)\}=\emptyset$, so that $\tau_{t}^{x}=\inf\{t\leq s\leq T|v(s,F_{s}^{t,x})=\varphi(s,F_{s}^{t,x})\}=T$}. Now since $\tau_{t}^{x}$ is the smallest optimal stopping time for (7), as per Theorem 3.3 (iii), we conclude that it is unique.
@@ 4.5. Equivalence of the Optimal Stopping Problems
If $L(t)<0$ for all $t\in[0,T)$, v(t,x)>h(t,x){v(t,x)>h(t,x)} for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, where $h(t,x)=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right]$}.
@@
To show that $\mathcal{S}\supseteq\widetilde{\mathcal{S}}$, we show $\mathcal{C}\subseteq\widetilde{\mathcal{C}}$. Fix $(t,x)\in\mathcal{C}$, so that $v(t,x)>\varphi(t,x)$}. Since $L(t)<0$ for all $t\in[0,T)$, we have by Lemma 4.5 that v(t,x)>h(t,x){v(t,x)>h(t,x)} for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$. Thus, $v(t,x)>\max(\varphi(t,x),h(t,x))$, which confirms that $(t,x)\in\widetilde{\mathcal{C}}$. ∎
@@
[…]
@@
To provide some intuition for this result, we remark that the condition $L(t)<0$ is associated with a non-empty surrender region (see Proposition 4.5); at any point during the life of the contract, there is always a possibility that early surrender will become optimal in the future. Then, the value of the VA contract is strictly above that of the EPV of the maturity benefit, and if surrender occurs it will be due to the contract value being equal to the surrender benefit. When this is the case, all three stopping times coincide (see the discussion after the proof of Lemma 3.3).
@@ 5. Numerical Examples
5.1. Time-dependent fee function
In this section, we present two simple examples to highlight specific features that can be observed in the optimal stopping region associated with variable annuity pricing: its disconnectedness and the discontinuity of its boundary.
@@ 5.1. Time-dependent fee function
$c_{1}(t)$ $=\frac{1}{100}\left(0.00889t^{2}-0.1330t+1\right),\qquad c_{2}(t)$ $=\frac{1}{100}\left(-0.0333t+1\right),$
for $0\leq t\leq T$, with $T=15$. The surrender charge function is set to $g(t)=e^{-0.0055(T-t)}$, a commonly used form in the variable annuity literature. The resulting surrender charge at inception of the contract is close to $8\%$, and it goes down to less than $0.5\%$ at $t=14$, one year before maturity.
The fee functions were chosen to illustrate some of the results obtained in Section 4. Function $c_{1}$ shows the impact of a change in the sign of $L(t)$; it is negative on $0,5.17]\cup[9.79,15]$ and positive elsewhere. Function $c_{2}$ was chosen such that $L(t)>0$ for $t\in(13.5,T]$ for $c_{2}$, thus eliminating the risk-neutral surrender incentive in the last 1.5 year of the contract. In both cases, the annual fee rates vary between $0.5\%$ and $1.0\%$. In practice, the fee is typically set as a constant. Numerical examples in this setting are provided, for example, in Bernard et al. (2014b), MacKay (2014) and MacKay et al. (2017).
@@
Figure 1 shows the continuation regions (in red) associated with the optimal stopping problem (7) for the two fee functions $c_{1}$ and $c_{2}$. The surrender regions (in white) are empty between years $5$ and $10$ for the fee function $c_{1}$, and between years $13.5$ and $15$ for $c_{2}$. This corresponds to the time intervals during which $L(t)>0$, illustrating a result of Proposition 4.6 (ii). Thus, the optimal surrender boundary is discontinuous when the fee is given by $c_{1}$. Other examples of the influence of the fee function on the surrender region exist, for example, in MacKay et al. (2017), Figures 4 and 5. However, disconnected sets for the surrender region have not been observed in prior numerical work on similar problems, see among others Bernard et al. (2014b), MacKay et al. (2017), Kang and Ziveyi (2018), and MacKay et al. (2023).
Another important feature highlighted in the second example (with fee function $c_{2}$) is that $\lim_{t\rightarrow T}b(t)=G$ does not hold. This is because the surrender region is empty for $t>13.5$, just before maturity (since $L(t)>0$). Moreover, in this particular case, since $\mathcal{S}\cap([13.5,T)\times\mathbb{R}_{+})=\emptyset$, we have $v(t,x)=h(t,x)$ for $(t,x)\in[13.5,T)$, so that $\widetilde{\mathcal{S}}\cap([13.5,T)\times\mathbb{R}_{+})=[13.5,T)\times\mathbb{R}_{+}$, providing another example in which the optimal stopping problems with the discontinuous and the continuous reward functions have different surrender (and continuation) regions.
Finally, we note that in our figures, the surrender region becomes empty slightly before $t=5$ for the fee function $c_{1}$, and before $t=13.5$ for $c_{2}$. This is likely due to the $y$-axis stopping at 500; it may be that there exists a surrender region at $t=5$ for account values that are (much) higher than 500, and that the surrender boundary is continuous in a general sense, “reaching infinity” continuously when the sign of $L$ changes. Thus, we conjecture that $L(t)\leq 0$ if and only if $\mathcal{S}_{t}$ is non-empty, but this is not something we have been able to show.
@@
5.2. State-dependent fee function
Next, to show the impact of a fee function that depends on the value of the underlying account, we define
$c_{3}(t,x)=\frac{0.12e^{150-x}}{1+e^{150-x}}.$
This function was chosen to resemble the discontinuous fee studied in MacKay et al. (2017) while satisfying Assumption 4.1. The resulting fee stays very close to 0.012 until the investment account $F_{t}$ reaches values around 150, at which point it drops quickly to 0. The idea behind such a state-dependent fee is to reduce the surrender incentive by eliminating the fee when the value of the maturity guarantee is low. Following MacKay et al. (2017), we also use $g(t,x)=1-0.05\left(1-\frac{t}{T}\right)^{3}$, $F_{0}=G=100$, $T=10$, $\sigma=0.165$ and $r=0.03$. The resulting optimal surrender region is presented in Figure 2.
Figure 2. Continuation region (in red) of the optimal stopping problem (7). The value of the variable annuity contract is approximated using the continuous-time Markov chain approximation described in MacKay et al. (2023). Market and VA parameters are $r=0.03$, $\sigma=0.165$, $F_{0}=G=100$ and $T=10$.
This specific state-dependent fee removes the optimal surrender incentive on $0\leq t\leq 5$ (that is, $\mathcal{S}\cap[0,5]=\emptyset$). Indeed, whenever the $t$-section of the surrender region is non-empty, it has the form $(b_{1}(t),b_{2}(t))$, for $b_{1}(t)$, $b_{2}(t)<\infty$. The surrender region is not of the so-called “threshold” type; if the account value increases quickly at the beginning of the contract and remains high until maturity, it may never be optimal to surrender. However, it is optimal to do so for some values of $x$ on $t\in(5,10]$. This is due to the vanishing fee as $F_{t}$ increases.
@@ 6. Conclusion
In particular, the continuous reward representation allows us to directly apply results from optimal stopping theory to obtain the continuity of the value function. We also prove its convexity in the state variable, its Lipschitz property in $x$ and characterize its continuity in $t$. Then, we derive an integral expression for the early surrender premium, generalizing the results of Bernard et al. (2014a) to general time-dependent fee and surrender charge functions. This second representation of the value function is also known as the early exercise premium representation in the American option pricing literature. From there, we develop a third representation for the value function in terms of the current surrender value and an integral expression that only takes value in the continuation region. We call this third representation the continuation premium representation.
@@
[…]
    $\quad\left.\left.-ye^{(r-\sigma^{2}/2)(\tau^{x}_{t}-t)-\int_{0}^{\tau_{t}^{x}-t}c(t+s)\,\mathrm{d}s+\sigma W_{\tau^{x}_{t}-t}}\right|\right]$
@@
[…]
@@
[…]
    Fix $0\leq s\leq t\leq T$. Since $v$ is non-monotone in $t$ (see Remark 4.5), we consider two cases: $v(s,x)\geq v(t,x)$ and $v(s,x)<v(t,x)$ .
@@
    $\qquad\qquad+|\sqrt{t-s}|x\sigma e^{|r-\sigma^{2}/2|T+\int_{0}^{T}c(u)\,\mathrm{d}u}\mathbb{E}[M_{1}e^{\sigma\sqrt{T}M_{1}}]$
    $\quad=\widetilde{C}_{1}|t-s|+C_{2}|\sqrt{t-s}|,$
@@
    $\quad\leq C_{1}|t-s|+C_{2}\left|\sqrt{t-s}\right|$